You need to add a salt to a DNA solution. At the moment the DNA is dissolved in water, you need it to be in 0.2 M sodium acetate. The volume of the DNA solution is 8.8 ml and you have a stock 2 M sodium acetate. 

How much sodium acetate do you need to add to the DNA solution to bring the final concentration to 0.2 M sodium acetate?


1
Expert's answer
2020-07-23T05:32:34-0400

0.2M is required (0.2 M = 0.0002 mol/ml). c=n/v where n= moles and v =volume. The stock solution has a conc. of 2M( = 0.002 mol/ml). Let x = ml of stock added.

Now n = cv. (n = 0.002x, v = 8.8 + x). Required Conc. = 0.0002 mol/ml

So, 0.0002 = n/v = 0.002x/(8.8 + x) . Solve the equation for x.

Algebraically x = 0.978 ml of stock added.